What volume of 8.50 M H3PO4 is required to prepare 204 mL of a 0.750 M solution?
1
Expert's answer
2020-11-26T12:20:19-0500

Solution:

In order to get the answer, you have to use the dilution equation:

CM1 × V1 = CM2 × V2

where:

CM1 is the initial concentration (8.50 M)

V1 is the initial volume (unknown)

CM2 is the final concentration (0.750 M)

V2 is the final volume (204 mL)


We know CM1, CM2, and V2.

Thus:

V1 = CM2 × V2 / CM1

V1 = (0.750 M × 204 mL) / 8.50 M = 18 mL

V1 = 18 mL.


Answer: 18 mL of H3PO4.